OpenStudy (anonymous):

How to Graph 4x-12=0. Can you explain step by step

OpenStudy (anonymous):

mathway.com

OpenStudy (mathmale):

After you've gotten some experience with this type of equation, it'll be easy for you. Since 4x-12=0 has only one variable, we can solve for that variable and get a result that is a constant: 4x=12 => x = 3. Mark x=3 on the horizontal axis and draw a vertical line through it. That's the graph of 4x-12=0.

OpenStudy (campbell_st):

well its interesting... in that all you can do is put a point on the number line at the solution. to graph on the number plane you would need 2 variables... x and y so the equation might be y = 4x - 12 you have 4x - 12 = 0 which has a solution of x = 3... draw a number line and put a dot on 3... hope it makes sense

OpenStudy (campbell_st):

the problem is 4x - 12 = 0 is really asking where do y = 0 and y = 4x -12 intersect.... perhaps its a typo and you left out the y.